Lead(II) sulfate formed in the reaction of 145 mL of 0.123 M lead(II) nitrate and excess sodium sulfate.


c. Calculate the mass of lead(II) sulfate formed in the reaction



1
Expert's answer
2022-03-16T05:22:05-0400

Solution:

lead(II) nitrate = Pb(NO3)2

sodium sulfate = Na2SO4

lead(II) sulfate = PbSO4


Calculate the moles of Pb(NO3)2:

Moles of Pb(NO3)2 = Molarity of Pb(NO3)2 × Volume of Pb(NO3)2

Moles of Pb(NO3)2 = (0.123 M) × (0.145 L) = 0.017835 mol


Balanced chemical equation:

Pb(NO3)2 + Na2SO4 → PbSO4 + 2NaNO3

According to stoichiometry:

1 mol of Pb(NO3)2 produces 1 mol of PbSO4

Thus, 0.017835 ol of Pb(NO3)2 produces:

[0.017835 mol Pb(NO3)2] × [1 mol PbSO4 / 1 mol Pb(NO3)2] = 0.017835 mol PbSO4


Calculate the mass of PbSO4:

The molar mass of PbSO4 is 303.26 g/mol

Therefore,

Mass of PbSO4 = (0.017835 mol PbSO4) × (303.26 g PbSO4 / 1 mol PbSO4) = 5.41 g PbSO4

Mass of PbSO4 = 5.41 g PbSO4


Answer: 5.41 grams of lead(II) sulfate (PbSO4) formed in the reaction